OpenStudy (anonymous):

Evaluate : cos2 32°+ cos2 58°/sin 2 59°+ sin 2 31° Evaluate : cos2 32°+ cos2 58°/sin 2 59°+ sin 2 31° @Mathematics

OpenStudy (yash2651995):

is it square or twice?

OpenStudy (anonymous):

TWICE

OpenStudy (yash2651995):

you sure??

OpenStudy (anonymous):

Put it in your calculator.

OpenStudy (anonymous):

i am sure

OpenStudy (yash2651995):

Evaluate : cos 64°+ cos 116°/sin 118°+ sin 62° cos 116 =cos (180-64) = sin 64 cos 118 =cos (180-62) = sin 62

OpenStudy (yash2651995):

Evaluate : cos 64°+ sin 64°/cos 62°+ sin 62°